OpenStudy (anonymous):

the perimeter of a rectangle is 118 inches. the length exceeds the width by 27 iches. find the length and the width.

OpenStudy (anonymous):

let lenght be 27 + x and breadth be x 27 +x +x = 118 2x = 118-27 2x= 91 x = 91/2

OpenStudy (anonymous):

if P=2W +2L W+27 = L since rectangle have 2 width sides and 2 length sides we get: (w +27) + (w +27) [these are the length] + w + w [width] w^4 + 54 = 118 w=16 thus, L = 43 and W = 16 2L + 2W or (86 + 32) = 118

OpenStudy (anonymous):

@ Aryan: not sure if thats the formula to find the area of the rectangle possibly.

OpenStudy (anonymous):

sorry i did it wrong

OpenStudy (anonymous):

iz cool bro. i just looked for a sec and thought " is that...huh....ok hold right hurr...lolwut?

OpenStudy (anonymous):

let l be 27+x and b be x 2(27+x +x) = 118 27 + x + x = 118/2 27 + 2x = 59 2x = 59 - 27 2x = 32 x = 32/2 x = 16 l = 27 + 16 = 43 inches b = 16 inches